how many sodium ions are present in 400cm3 of 0.5mol/dm^3 of sodium trioxocarbonate 5 solution


1
Expert's answer
2021-03-15T09:45:46-0400

400cm3= 0.4dm3

Moles of sodium ions=0.2moles

If 1mole =6.022×10^23 ions

Then;

"0.2\u00d7 6.022\u00d710^23" =1.2×10^23 ions